# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v2

@@ 10.6 Overlapping HR Protections
Let’s start with a key observation. In India, when institutions announce positions for recruitment or admissions, they categorize these positions based on vertical categories (including open) and horizontal protections (or lack thereof). For example, in a setting with four VR-protected categories—SC, ST, OBC, and EWS—and two HR-protected groups—PwD and W—specific numbers of positions are designated for pairs such as SC-W or Open-PwD. In practice, this means that when an individual belonging to multiple HR-protected groups is awarded a position, only one of their HR protections is honored. Thus, if a woman with disabilities is awarded a position, she fulfills either the HR protection for W or for PwD, but not both.118118118This convention is referred to as one-to-one HR matching in Sönmez and Yenmez (2022). See Sönmez and Yenmez (2020) for an analysis of the alternative one-to-all HR matching convention, in which an individual fulfills all their HR protections.
@@
In the current setting, however, challenging an individual’s assignment may require adjustments to one or more other assignments. Specifically, although the challenge is directed at a particular individual, the correction may also involve others. In Example 8, for instance, Chandra’s potential objection to Debraj’s assignment in the first scenario would also entail a change in Amita’s assignment from Open-W to Open-PwD. While, in my opinion, such a challenge aligns with Saurav Yadav vs. State of U.P. (2020),119119119As the basis for my opinion, see Mamta Bisht vs. State of Uttaranchal (2005) at the Uttarakhand High Court, where the plaintiff successfully challenged an individual’s assignment in the open category indirectly. She argued that a third party who had been awarded an SC position should instead be assigned the open category position, thereby allowing the SC position to be reassigned to the plaintiff. as an economist, I cannot assert that this point is clear-cut under reservation jurisprudence.
@@ 10.6.1 Smart Reserves and Meritorious Horizontal Choice Rule
This idea extends naturally regardless of the specifics of the HR-protected groups or their overlapping memberships. First, we allocate the HR-protected positions in a “greedy” manner to the highest merit-score individuals, maximizing the total number of these positions that are honored.120120120Technically, this requires solving a maximum matching problem in a bipartite graph. In an unweighted bipartite graph, the objective is to find a matching with maximum cardinality. The Hopcroft-Karp algorithm (Hopcroft and Karp, 1973) solves this problem efficiently, and even faster algorithms exist for specific classes of bipartite graphs. Then, we allocate the remaining positions in the category to the highest merit-score eligible individuals who have not yet received a position. Introduced in Sönmez and Yenmez (2020), we refer to this choice rule as the meritorious horizontal choice rule.121121121The terminology meritorious horizontal choice rule was first introduced in Sönmez and Yenmez (2022). Originally, Sönmez and Yenmez (2020) referred to this choice rule as the horizontal envelope choice rule.
@@ 10.7 Allocation of Heterogenous Positions
The creation of such “artificial” property rights for specific categories of individuals based on flawed mechanics not only represents a misguided attempt to resolve the underlying issues with the SSD-based mechanism but also directly contradicts the axiom of no justified envy. Thus, the root cause of these legal challenges lies in the excessive reliance on the concept of “migration,” further exacerbated by the practice of awarding “artificial” property rights to arbitrary groups based on the tentative assignments of meritorious reserved candidates. This reliance on migration is an ill-equipped tool for addressing applications with heterogeneous positions and represents a methodological failure at the core of the legal inconsistencies and litigations in India.122122122A simple search on Indian Kanoon, a free search engine for Indian law, reveals that as of November 2024, there have been 959 cases related to the “migration of meritorious reserved candidates” at the Supreme Court or various high courts.
@@ 11.1 The Toughest Triage
Upon exploring with Parag Pathak and Utku Ünver how these ethical principles were operationalized for allocation of ventilators and ICUs, we made an important discovery, one that made it clear that it is time to put our expertise into good use (Pathak et al., 2020c). Little did we know that, for nearly a whole year from that time on, the three of us would end up setting aside all unessential activities and devote all our time and energy to contribute to the global fight to combat COVID-19.
@@ 11.2 Priority Point System: Prevalence, Challenges, and Constraints
As their recommended allocation rule for crisis rationing of ventilators and ICUs, state guidelines in the U.S. almost uniformly suggest some form of a priority point system. In these guidelines, a patient’s point score is most often determined by a measure of mortality risk called the Sequential Organ Failure Assessment (SOFA) score.123123123This procedure can be thought as a simple serial dictatorship induced by the SOFA score. I refer to this specific priority point system as the SOFA system.124124124SOFA scoring system is first introduced in Vincent et al. (1996) on behalf of the Working Group on Sepsis-Related Problems of the European Society of Intensive Care Medicine. Under this system, each of six organ groups lungs, liver, brain, kidneys, blood clotting and blood pressure is assigned a score of 1 to 4, with higher scores for more severely failed organs. The total across the six organ groups determine the SOFA score of the patient, which then determines her priority for a ventilator or an ICU. In recent years, however, many expressed concerns that the SOFA system discriminates against certain groups. See, for example, Tolchin et al. (2021) for one such study.
@@
For example, at the onset of COVID-19, the leading allocation system that accommodated multiple ethical principles for the allocation of ventilators was a priority point system by White et al. (2009). Through a modification of the SOFA system, their system incorporated three ethical principles as follows: It utilized a coarsening of patient SOFA scores to accommodate the principle of saving most lives and added points to address two additional principles, saving the most years of life and the life-cycle principle.125125125The life-cycle principle asserts that younger patients, who have had the least opportunity to live through various stages of life, should be given higher priority for the scarce medical resource.
@@ 11.3 Pandemic Reserve System: A Minimalist Resolution
In real-world settings, reserve systems have been used for decades as a prominent tool for reaching compromises between groups with competing objectives. These settings include ensuring equitable representation of minorities and women in legislative assemblies (Htun, 2004; Dahlerup, 2007; Krook, 2009; Hughes, 2011), implementing affirmative action for allocating government positions and public school seats (Correa et al., 2019; Dur et al., 2020; Aygün and Bó, 2021; Sönmez and Yenmez, 2022), balancing school seat priorities between neighborhood populations and the broader community (Dur et al., 2018b), and allocating immigration visas in the U.S. (Pathak et al., 2020a). However, given the diverse allocation criteria for different medical resources across numerous jurisdictions, the urgent need to allocate scarce medical resources globally during the pandemic required the examination of a more generalized class of reserve systems (Pathak et al., 2024).
@@
Collaborating with Bumin Yenmez and working around the clock to make a timely contribution to the global fight against COVID-19, Pathak, Ünver, and I completed and circulated the first version of our paper Pathak et al. (2020d) in April 2020, just two weeks after beginning this critical resource allocation project. The final version was later published in Pathak et al. (2024).
@@
We observed a prime example of this issue in Section 9.2, where the flawed implementation of walk zone priorities at Boston Public Schools from 2005 to 2013 led to unintended outcomes. Similarly, in Section 10, we examine this problem in the context of affirmative action policies and court rulings in India. This subtlety has also affected the implementation of H-1B visa assignments in the U.S. (Pathak et al., 2020a). To prevent confusion among healthcare officials implementing our proposed reserve system for allocating vital medical resources during the COVID-19 pandemic, we focused our formal analysis in Pathak et al. (2020d, 2024) on addressing this subtle aspect of the problem.
As a side note, while Pathak et al. (2020d) marked the first time we used minimalist market design in real-time to offer a resolution to an important crisis, it was not the last. When the contested 103rd Constitutional (Amendment) Act, 2019 was deliberated in the hearings of a recent Supreme Court judgment Janhit Abhiyan vs. Union of India (2022) in November 2022, in Sönmez and Ünver (2022), we offered a real-time resolution before a divisive decision was declared by the court in a split 3-2 vote, with strong protests from the two dissenting justices.126126126Indeed, about a week before the split decision, in Deshpande and Sönmez (2023), I collaborated with Ashwini Deshpande, a prominent expert on the Indian reservation system, to publish an op-ed in The Hindu. Together, we cautioned Indian society on the subtleties of the case. These experiences highlight the utility of minimalist market design—a highly structured framework based on identifying and remedying root causes of failures—in providing timely resolutions for important crises when they are most useful for society.
@@ 11.4 Teaming Up with Three Leaders in Bioethics
Timing is often everything in market design. A mere two days after Pathak et al. (2020d) was circulated, the following headline from a lead story in the Boston Globe caught our attention (Kowalczyk, 2020):
@@
The system recommended in the Original Guidance was basically a simplified version of the priority point system of White et al. (2009), with only two ethical principles instead of three. Naturally, we wondered how the Boston Globe could possibly make such a major error in its coverage. Subsequently, we observed that parts of the Original Guidance used ambiguous language, mentioning possible preference given to medical personnel who are vital to treating others, even though the final recommendation did not reflect this consideration.127127127The ambiguous language in the Original Guidance was later removed in a Revised Guidance issued on April 20, 2020, possibly due to an inquiry by our team about this inconsistency to Robert Truog, who was a member of the working group mentioned in the Boston Globe story.
@@
Based on this hunch, later that evening we sent an email to Robert Truog, who was mentioned in the Boston Globe story as a member of the working group. Sharing our ideas with him was especially important—not only was he the Director of the Center for Bioethics at Harvard Medical School, but he was also the first author of Truog et al. (2020), one of the two articles that initially sparked our interest in the topic. In our email, we attached our paper Pathak et al. (2020d), mentioned that we were inspired by the NEJM symposium that included his paper, and expressed our strong interest in exchanging ideas with him in a Zoom meeting soon.
@@ 11.5.2 NASEM Framework for Equitable Vaccine Allocation
Anticipating this need and in collaboration with Schmidt, we circulated an NBER working paper Pathak et al. (2020b) weeks earlier, illustrating how a tiered priority system can be easily modified into a reserve system as a minimalist refinement by incorporating equity through an index of social vulnerability. Schmidt brought this exercise in minimalist market design to the committee’s attention as a possible mechanism to embed equity into their framework.
In October 2020, NASEM announced its Framework for Equitable Allocation of COVID-19 Vaccine (NASEM, 2020a). Following the recommendation in Persad et al. (2020) and using the exact formulation from Pathak et al. (2020b), the NASEM framework formally recommended a 10 percent reserve for people from hard-hit areas:
@@
Categories, their share of COVID-19 vaccines, and factors that affect category-specific priorities were specified by the health districts as follows:128128128Category-specific priorities are only partially specified in RichmondaCity (2021).
@@
Notably, the Richmond and Henrico reserve system utilized the full generality of the model presented in Pathak et al. (2020d, 2024). Specifically, while categories 2 and 3 used the same baseline priority ranking (subject to eligibility), category 4 employed a related but different priority ranking.
@@ 12.2 A Biological Incentive Mechanism for Liver Exchange
Another crucial statistic reveals that, in much of the world, donor mortality and morbidity rates are much higher for right-lobe hepatectomy than for left-lobe hepatectomy (Lee, 2010; Mishra et al., 2018).129129129Lee (2010) reports mortality rates of 0.4-0.5% for right hepatectomy and 0.1% for left hepatectomy. Mishra et al. (2018) reports morbidity rates of 28% for right hepatectomy and 7.5% for left hepatectomy. Consequently, while the left lobe is frequently deemed unsuitable for transplantation, it represents a much safer option for the donor. This observation forms the basis for a “biological” incentive that may compel blood-type-compatible patient-donor pairs to participate in liver exchange.
Consider a scenario involving a blood type A patient with a petite blood type O donor. The patient can only receive the larger right lobe of the donor. In a typical liver exchange pool, there are often numerous blood type O patients who are incompatible with their corresponding blood type A donors.130130130In matching theory, we refer such pairs as “under-demanded” (Roth et al., 2005b, 2007). Frequently, some of them—pediatric or slender individuals—require smaller liver grafts. While the first patient is compatible with their donor, unlike the case of a kidney transplant, a donor exchange between the two pairs becomes a visibly appealing option. This is due to the fact that the first donor can undergo a much less risky left-lobe hepatectomy through liver exchange, as opposed to the considerably higher-risk direct donation to their own patient.
@@ 12.6 Banu Bedestenci Sönmez Liver Paired Exchange System
In June 2022, we received the initial dataset from Malatya and began running the system.131131131Before the deployment of the computerized system in June 2022, the Malatya team conducted 2-way exchanges on an emergency basis on two occasions: first in October 2021 and second in January 2022. Just days before my first visit to Turkey in three years, during the first week of July 2022, we surprised the Malatya team with an unexpected outcome that no one had anticipated, especially not at such an early stage.

### v3

@@ 10.6 Overlapping Horizontal Reservations
Let’s start with a key observation. In India, when institutions announce positions for recruitment or admissions, they categorize these positions based on vertical categories (including open) and horizontal protections (or lack thereof). For example, in a setting with four VR-protected categories—SC, ST, OBC, and EWS—and two HR-protected groups—PwD and W—specific numbers of positions are designated for pairs such as SC-W or Open-PwD. In practice, this means that when an individual belonging to multiple HR-protected groups is awarded a position, only one of their HR protections is honored. Thus, if a woman with disabilities is awarded a position, she fulfills either the HR protection for W or for PwD, but not both.122122122This convention is referred to as one-to-one HR matching in Sönmez and Yenmez (2022). See Sönmez and Yenmez (2020) for an analysis of the alternative one-to-all HR matching convention, in which an individual fulfills all their HR protections.
@@
In the current setting, however, challenging an individual’s assignment may require adjustments to one or more other assignments. Specifically, although the challenge is directed at a particular individual, the correction may also involve others. In Example 8, for instance, Chandra’s potential objection to Debraj’s assignment in the first scenario would also entail a change in Amita’s assignment from Open-W to Open-PwD. While, in my opinion, such a challenge aligns with Saurav Yadav vs. State of U.P. (2020),123123123As the basis for my opinion, see Mamta Bisht vs. State of Uttaranchal (2005) at the Uttarakhand High Court, where the plaintiff successfully challenged an individual’s assignment in the open category indirectly. She argued that a third party who had been awarded an SC position should instead be assigned the open category position, thereby allowing the SC position to be reassigned to the plaintiff. as an economist, I cannot assert that this point is clear-cut under reservation jurisprudence.
@@ 10.6.1 Smart Reserves and Meritorious Horizontal Choice Rule
This idea extends naturally regardless of the specifics of the HR-protected groups or their overlapping memberships. First, we allocate the HR-protected positions in a “greedy” manner to the highest merit-score individuals, maximizing the total number of these positions that are honored.124124124Technically, this requires solving a maximum matching problem in a bipartite graph. In an unweighted bipartite graph, the objective is to find a matching with maximum cardinality. The Hopcroft-Karp algorithm (Hopcroft and Karp, 1973) solves this problem efficiently, and even faster algorithms exist for specific classes of bipartite graphs. Then, we allocate the remaining positions in the category to the highest merit-score eligible individuals who have not yet received a position. Introduced in Sönmez and Yenmez (2020), we refer to this choice rule as the meritorious horizontal choice rule.125125125The terminology meritorious horizontal choice rule was first introduced in Sönmez and Yenmez (2022). Originally, Sönmez and Yenmez (2020) referred to this choice rule as the horizontal envelope choice rule.
@@ 10.7 Allocation of Heterogenous Positions
The creation of such “artificial” property rights for specific categories of individuals based on flawed mechanics not only represents a misguided attempt to resolve the underlying issues with the SSD-based mechanism but also directly contradicts the axiom of no justified envy. Thus, the root cause of these legal challenges lies in the excessive reliance on the concept of “migration,” further exacerbated by the practice of awarding “artificial” property rights to arbitrary groups based on the tentative assignments of meritorious reserved candidates. This reliance on migration is an ill-equipped tool for addressing applications with heterogeneous positions and represents a methodological failure at the core of the legal inconsistencies and litigations in India.126126126A simple search on Indian Kanoon, a free search engine for Indian law, reveals that as of November 2024, there have been 959 cases related to the “migration of meritorious reserved candidates” at the Supreme Court or various high courts.
@@ 11.1 The Toughest Triage
Upon exploring with Parag Pathak and Utku Ünver how these ethical principles were operationalized for allocation of ventilators and ICUs, we made an important discovery, one that made it clear that it is time to put our expertise into good use (Pathak et al., 2020b). Little did we know that, for nearly a whole year from that time on, the three of us would end up setting aside all unessential activities and devote all our time and energy to contribute to the global fight to combat COVID-19.
@@ 11.2 Priority Point System: Prevalence, Challenges, and Constraints
As their recommended allocation rule for crisis rationing of ventilators and ICUs, state guidelines in the U.S. almost uniformly suggest some form of a priority point system. In these guidelines, a patient’s point score is most often determined by a measure of mortality risk called the Sequential Organ Failure Assessment (SOFA) score.127127127This procedure can be thought as a simple serial dictatorship induced by the SOFA score. I refer to this specific priority point system as the SOFA system.128128128SOFA scoring system is first introduced in Vincent et al. (1996) on behalf of the Working Group on Sepsis-Related Problems of the European Society of Intensive Care Medicine. Under this system, each of six organ groups lungs, liver, brain, kidneys, blood clotting and blood pressure is assigned a score of 1 to 4, with higher scores for more severely failed organs. The total across the six organ groups determine the SOFA score of the patient, which then determines her priority for a ventilator or an ICU. In recent years, however, many expressed concerns that the SOFA system discriminates against certain groups. See, for example, Tolchin et al. (2021) for one such study.
@@
For example, at the onset of COVID-19, the leading allocation system that accommodated multiple ethical principles for the allocation of ventilators was a priority point system by White et al. (2009). Through a modification of the SOFA system, their system incorporated three ethical principles as follows: It utilized a coarsening of patient SOFA scores to accommodate the principle of saving most lives and added points to address two additional principles, saving the most years of life and the life-cycle principle.129129129The life-cycle principle asserts that younger patients, who have had the least opportunity to live through various stages of life, should be given higher priority for the scarce medical resource.
@@ 11.3 Pandemic Reserve System: A Minimalist Resolution
In real-world settings, reserve systems have been used for decades as a prominent tool for reaching compromises between groups with competing objectives. These settings include ensuring equitable representation of minorities and women in legislative assemblies (Htun, 2004; Dahlerup, 2007; Krook, 2009; Hughes, 2011), implementing affirmative action for allocating government positions and public school seats (Correa et al., 2019; Dur et al., 2020; Aygün and Bó, 2021; Sönmez and Yenmez, 2022), balancing school seat priorities between neighborhood populations and the broader community (Dur et al., 2018b), and allocating immigration visas in the U.S. (Pathak et al., 2025). However, given the diverse allocation criteria for different medical resources across numerous jurisdictions, the urgent need to allocate scarce medical resources globally during the pandemic required the examination of a more generalized class of reserve systems (Pathak et al., 2024).
@@
Collaborating with Bumin Yenmez and working around the clock to make a timely contribution to the global fight against COVID-19, Pathak, Ünver, and I completed and circulated the first version of our paper Pathak et al. (2020c) in April 2020, just two weeks after beginning this critical resource allocation project. The final version was later published in Pathak et al. (2024).
@@
We observed a prime example of this issue in Section 9.2, where the flawed implementation of walk zone priorities at Boston Public Schools from 2005 to 2013 led to unintended outcomes. Similarly, in Section 10, we examine this problem in the context of affirmative action policies and court rulings in India. This subtlety has also affected the implementation of H-1B visa assignments in the U.S. (Pathak et al., 2025). To prevent confusion among healthcare officials implementing our proposed reserve system for allocating vital medical resources during the COVID-19 pandemic, we focused our formal analysis in Pathak et al. (2020c, 2024) on addressing this subtle aspect of the problem.
As a side note, while Pathak et al. (2020c) marked the first time we used minimalist market design in real-time to offer a resolution to an important crisis, it was not the last. When the contested 103rd Constitutional (Amendment) Act, 2019 was deliberated in the hearings of a recent Supreme Court judgment Janhit Abhiyan vs. Union of India (2022) in November 2022, in Sönmez and Ünver (2022), we offered a real-time resolution before a divisive decision was declared by the court in a split 3-2 vote, with strong protests from the two dissenting justices.130130130Indeed, about a week before the split decision, in Deshpande and Sönmez (2023), I collaborated with Ashwini Deshpande, a prominent expert on the Indian reservation system, to publish an op-ed in The Hindu. Together, we cautioned Indian society on the subtleties of the case. These experiences highlight the utility of minimalist market design—a highly structured framework based on identifying and remedying root causes of failures—in providing timely resolutions for important crises when they are most useful for society.
@@ 11.4 Teaming Up with Three Leaders in Bioethics
Timing is often everything in market design. A mere two days after Pathak et al. (2020c) was circulated, the following headline from a lead story in the Boston Globe caught our attention (Kowalczyk, 2020):
@@
The system recommended in the Original Guidance was basically a simplified version of the priority point system of White et al. (2009), with only two ethical principles instead of three. Naturally, we wondered how the Boston Globe could possibly make such a major error in its coverage. Subsequently, we observed that parts of the Original Guidance used ambiguous language, mentioning possible preference given to medical personnel who are vital to treating others, even though the final recommendation did not reflect this consideration.131131131The ambiguous language in the Original Guidance was later removed in a Revised Guidance issued on April 20, 2020, possibly due to an inquiry by our team about this inconsistency to Robert Truog, who was a member of the working group mentioned in the Boston Globe story.
@@
Based on this hunch, later that evening we sent an email to Robert Truog, who was mentioned in the Boston Globe story as a member of the working group. Sharing our ideas with him was especially important—not only was he the Director of the Center for Bioethics at Harvard Medical School, but he was also the first author of Truog et al. (2020), one of the two articles that initially sparked our interest in the topic. In our email, we attached our paper Pathak et al. (2020c), mentioned that we were inspired by the NEJM symposium that included his paper, and expressed our strong interest in exchanging ideas with him in a Zoom meeting soon.
@@ 11.5.2 NASEM Framework for Equitable Vaccine Allocation
Anticipating this need and in collaboration with Schmidt, we circulated an NBER working paper Pathak et al. (2020a) weeks earlier, illustrating how a tiered priority system can be easily modified into a reserve system as a minimalist refinement by incorporating equity through an index of social vulnerability. Schmidt brought this exercise in minimalist market design to the committee’s attention as a possible mechanism to embed equity into their framework.
In October 2020, NASEM announced its Framework for Equitable Allocation of COVID-19 Vaccine (NASEM, 2020a). Following the recommendation in Persad et al. (2020) and using the exact formulation from Pathak et al. (2020a), the NASEM framework formally recommended a 10 percent reserve for people from hard-hit areas:
@@
Categories, their share of COVID-19 vaccines, and factors that affect category-specific priorities were specified by the health districts as follows:132132132Category-specific priorities are only partially specified in RichmondaCity (2021).
@@
Notably, the Richmond and Henrico reserve system utilized the full generality of the model presented in Pathak et al. (2020c, 2024). Specifically, while categories 2 and 3 used the same baseline priority ranking (subject to eligibility), category 4 employed a related but different priority ranking.
@@ 12.2 A Biological Incentive Mechanism for Liver Exchange
Another crucial statistic reveals that, in much of the world, donor mortality and morbidity rates are much higher for right-lobe hepatectomy than for left-lobe hepatectomy (Lee, 2010; Mishra et al., 2018).133133133Lee (2010) reports mortality rates of 0.4-0.5% for right hepatectomy and 0.1% for left hepatectomy. Mishra et al. (2018) reports morbidity rates of 28% for right hepatectomy and 7.5% for left hepatectomy. Consequently, while the left lobe is frequently deemed unsuitable for transplantation, it represents a much safer option for the donor. This observation forms the basis for a “biological” incentive that may compel blood-type-compatible patient-donor pairs to participate in liver exchange.
Consider a scenario involving a blood type A patient with a petite blood type O donor. The patient can only receive the larger right lobe of the donor. In a typical liver exchange pool, there are often numerous blood type O patients who are incompatible with their corresponding blood type A donors.134134134In matching theory, we refer such pairs as “under-demanded” (Roth et al., 2005b, 2007). Frequently, some of them—pediatric or slender individuals—require smaller liver grafts. While the first patient is compatible with their donor, unlike the case of a kidney transplant, a donor exchange between the two pairs becomes a visibly appealing option. This is due to the fact that the first donor can undergo a much less risky left-lobe hepatectomy through liver exchange, as opposed to the considerably higher-risk direct donation to their own patient.
@@ 12.6 Banu Bedestenci Sönmez Liver Paired Exchange System
In June 2022, we received the initial dataset from Malatya and began running the system.135135135Before the deployment of the computerized system in June 2022, the Malatya team conducted 2-way exchanges on an emergency basis on two occasions: first in October 2021 and second in January 2022. Just days before my first visit to Turkey in three years, during the first week of July 2022, we surprised the Malatya team with an unexpected outcome that no one had anticipated, especially not at such an early stage.